40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Message ID: 251     Entry time: Tue Aug 3 11:22:04 2010
Author: Megan 
Type: DailyProgress 
Category:  
Subject: PSL Table Updates 

Both the analyzer cavity and reference cavity are able to be locked, and were simultaneously locked for a while yesterday . However, the temperature control loop is not working because the computer is not seeing the temperature data, so it didn't stay locked . But now we know it's possible !

The mirrors, beam splitter, and photodetectors are set up and aligned to measure the beat of the two signals and I am currently working on mode matching to get the beams the same size at the photodetector.

FSS_RCTRANSPD and ACAV_TRANSPD have been calibrated:

FSS_RCTRANSPD = 1.98 mW/V

ACAV_TRANSPD = 2.18 mW/V

ELOG V3.1.3-